OpenStudy (anonymous):

Solve: a/4-9=5 A. -16 B. -56 C. 56 D. 16

terenzreignz (terenzreignz):

\[\Large \frac{a}4 - 9 = 5\] This right?

OpenStudy (anonymous):

a/4-9=5 is the equation. your equation came out a bit weird/messed up.

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

I'm pretty sure it involves that fraction though... anyway, linear equations in one variable make use of properties of equality...

terenzreignz (terenzreignz):

In particular, what happens if you add 9 to both sides of the equation?

OpenStudy (anonymous):

oh right! I forgot about that. it would be a/4=14

terenzreignz (terenzreignz):

That's right :D And then... you multiply both sides by 4... (to get rid of the denominator of a) Leaving you with...?

OpenStudy (anonymous):

and then do I multiply both sides by 4 or divide?

OpenStudy (anonymous):

ok so i would get a=56

OpenStudy (anonymous):

right?

terenzreignz (terenzreignz):

Correct :D

OpenStudy (anonymous):

ok, thank you very much! :)

terenzreignz (terenzreignz):

No problem :)